OpenStudy (anonymous):

simplify by removing factors of 1 45q^10/ 81q^12

OpenStudy (luffingsails):

Factors of 1? That isn't making sense. But here is what I would do. \[\frac{45q^{10}}{81q^{12}}=\frac{5*9q^{10}}{9*9q{12}}=\frac{5q^{10}}{9q^{12}}\] \[\frac{5q^{10}}{9q^{12}}=\frac{5q^{10}}{9q^{10}q^{2}}=\frac{5}{9q^2}\]